# Discovering Global False Negatives On the Fly for Self-supervised Contrastive Learning

### v1

@@ 1 Introduction
In the absence of reliable labels to determine whether a pair of data is positive or negative, many methods resort to instance discrimination. To this end, positive pairs are defined as distinct augmented views of the anchor data, while negative pairs are generated by sampling from the whole dataset excluding the anchor data, irrespective of their semantics (Chen et al., 2020a; Yuan et al., 2022). However, negative pairs produced through this method lack reliability. Specifically, augmented views from images sharing similar semantic meanings are incorrectly deemed negative, leading to their embeddings being pushed apart. This inadvertently encourages the model to discard crucial semantic information. We term these undesirable negative pairs as false negatives (FN). Figure 1 shows some examples of false negatives seen during training on ImageNet, where the anchor images are different from their negative samples, but semantically similar. The presence of false negatives detrimentally impacts the representations learned through contrastive learning (Saunshi et al., 2019), with this effect becoming more pronounced in large-scale datasets featuring numerous semantic concepts (Chen et al., 2022).
@@ 2 Related Work
Early efforts in SSL formulated pretext tasks to enable models to learn representations from unlabeled data. Examples include predicting the relative offset between two patches within the same image (Doersch et al., 2016), solving jigsaw puzzles (Noroozi and Favaro, 2017), colorizing grayscale images (Zhang et al., 2016), and unsupervised deep clustering (Caron et al., 2019). However, these methods necessitate carefully crafted pretext tasks, which may not always apply to diverse domains, leading to a lack of generality.
@@ 3 CL with Global False Negative Identification
$\mathcal{L}_{\mathrm{GCL}}(\mathbf{w})=\mathbf{E}_{\mathbf{x}_{i}\in\mathcal{D},A,A^{\prime}\in\mathcal{P}}\left[-\tau\log\frac{\exp\left(\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{\prime})/\tau\right)}{\sum_{\mathbf{x}\in\mathcal{S}_{i}^{-}}\exp(\mathrm{sim}(\mathbf{z}_{i},E_{\mathbf{w}}(\mathbf{x}))/\tau)}\right],$ (1)
@@ 3.1 Learning Dynamic Per-Anchor Thresholds
Instead, we choose an optimization-based approach to automatically learn the per-anchor threshold $\lambda_{i}$ to select the top $\alpha$most similar negative data for the $i$-th anchor. To achieve this, we cast the problem of finding the $(1-\alpha)$-quantile of all similarity scores between $\mathbf{x}_{i}$ and all other samples, i.e., $R_{i}=\{\mathrm{sim}(E_{\mathbf{w}}(\mathbf{x}_{i}),E_{\mathbf{w}}(\mathbf{x}))\mid\mathbf{x}\in\mathcal{S}_{i}^{-}\}$ as the following optimization problem (Ogryczak and Tamir, 2003):
@@
$\min_{\mathbf{w}}\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ $=\frac{1}{n}\sum_{\mathbf{x}_{i}\in\mathcal{D}}\mathbf{E}_{A,A^{\prime}}\left[-\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{\prime})+\tau\log(|\widetilde{\mathcal{S}}_{i}^{-}|g(\mathbf{w},\lambda_{i};\mathbf{x}_{i},\widetilde{\mathcal{S}}_{i}^{-}))\right],$ (3)
@@
Note that $\min_{\mathbf{w}}\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ can be viewed as a stochastic bilevel optimization problem (Ghadimi and Wang, 2018) since the minimization of $\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ involves the solution $\bm{\lambda}$ to a lower level problem in (2). However, the problem in (3) is more challenging that most bilevel problems in the literature (Ghadimi and Wang, 2018; Ji et al., 2021) because the lower-level problem in (2) is non-smooth and non-strongly convex while the upper-level function $\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ is non-differentiable to $\bm{\lambda}$. To tackle this challenge, we just ignore the hypergradient of $\bm{\lambda}$ in terms of $\mathbf{w}$, which has been used in model-agnostic meta-learning (Finn et al., 2017).
@@ 3.2 GloFND for Unimodal CL
We propose an efficient algorithm called GloFND for dynamically discovering and eliminating the false negatives in contrastive learning. Our GloFND can be combined with previous contrastive learning algorithms, e.g., SogCLR (Yuan et al., 2022). In each iteration, SogCLR + GloFND first randomly samples a batch of data $\mathcal{B}\subset\mathcal{D}$ and data augmentations $A,A^{\prime}$. Then, it alternatively executes two steps: (i) updating the thresholds $\lambda_{i},i\in\mathcal{B}$; (ii) removing the identified false negatives from the loss function and updating the parameters $\mathbf{w}$ of the encoding network.
@@ 3.2.1 Updating the threshold $\lambda$
where $\mathbb{I}(\cdot)$ is the indicator function, $\Pi_{[-1,1]}[\cdot]$ denotes the projection onto the interval $[-1,1]$ due to that are similarity scores are in $[-1,1]$, and $\theta$ is the learning rate of $\lambda_{i}$. In this way, we keep track of a threshold $\lambda_{i}$ for detecting global false negatives in the whole dataset for each anchor $\mathbf{x}_{i}\in\mathcal{D}$.
@@ 3.2.2 Updating Encoder Network $E_{\mathbf{w}}$
Our approach GloFND can be extended to resolve the global false negative discovery in bimodal CL, e.g., CLIP (Radford et al., 2021). This can be achieved by learning a threshold for each instance for two modalities and following the same general procedure as for unimodal CL. We refer readers to Appendix B for more details.
@@ 4 Experiments
For unimodal and semi-supervised experiments, we compare GloFND with SogCLR (Yuan et al., 2022) and its FNC (Huynh et al., 2022) version. FNC computes the top $k$ for negative data within a mini-batch by utilizing a support set. The support set includes additional views for each image, and the similarity scores are averaged across these views. For a fair comparison, we set $k=\alpha|\mathcal{D}|$ and use a support set of size 1. For bimodal experiments, we compare with SogCLR, FastCLIP (Wei et al., 2024), and their FNC versions.
@@ 4.1 Unimodal and Semi-supervised Experiments
Dataset. We run our experiments on ImageNet100 (Wu et al., 2019), a subset of ImageNet with 100 randomly selected classes (about 128k images), and report scores on its official validation split. Additionally, we examine the transfer learning performance on Food-101 (Bossard et al., 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Stanford Cars (Krause et al., 2013), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Oxford-IIIT Pets (Parkhi et al., 2012), Caltech-101 (Li et al., 2022), and Oxford 102 Flowers (Nilsback and Zisserman, 2008).
Experiment Setup. Following previous work (Yuan et al., 2022), we pretrain ResNet-50 (He et al., 2015) with a 2-layer $128\times 128$ projection head on top of the backbone encoder. We pretrain for 200 epochs with a batch size of 128 and the same set of augmentations as in SogCLR. We use LARS optimizer (You et al., 2017) with square root learning rate scaling ($0.075\times sqrt(BatchSize)$) and cosine decay schedule without restart. For SogCLR, we set the temperature ($\tau$) to 0.1 and $\gamma=0.9$. We start using GloFND when we reach 70 epochs. We use $\alpha=0.01$, initialize $\lambda_{i}=1$, and learn it with Adam with a learning rate of $0.05$ ($\beta_{1}=0.9,\beta_{2}=0.98$) during the remaining epochs. For FNC, we set $\alpha=0.01$ and tune the starting epoch in $\{10,30,50,70,90,110,130\}$, choosing the value that achieves the best semi-supervised average performance.
@@
Results. We report false negative identification performance in Table 2. GloFND achieves significant improvements in false negative identification over FNC, with a 20.83% and 5.14% increase in mean precision and recall, leading to an F1-score of 53.10%, which is 16.68% higher than FNC. Tables 1 and 3 report linear evaluation and transfer learning results, respectively. Observe that simply removing the false negatives identified by FNC or GloFND improves both the semi-supervised and transfer learning performance of SogCLR. GloFND achieves greater improvements in both scenarios, achieving 1.04%-2.44% improvement in the semi-supervised scenario and an average 1.64% improvement in transfer learning, while increasing the per-epoch computation by only 2% (from 427 s to 435 s). More details can be found in Appendix I. Note these improvements are achieved by simply removing the false negatives identified by GloFND from the loss, while a more careful treatment can potentially improve the performance even further.
Table 1: Linear evaluation results in unimodal semi-supervised scenario. We train the linear classifiers with different percentages of randomly sampled labeled training data and present their top-1 accuracies (%) on the validation set. We include the overall average and, in parenthesis, its improvement WRT SogCLR baseline.
Method 100.0% 10.0% 1.0% 0.1% Average
SogCLR 76.55 (0.09) 72.24 (0.14) 62.92 (0.34) 34.94 (0.60) 61.66
+ FNC 77.12 (0.14) 72.89 (0.25) 64.29 (0.34) 36.11 (0.70) 62.60 (+0.94)
+ GloFND 77.59 (0.03) 73.36 (0.15) 65.09 (0.49) 37.38 (0.97) 63.36 (+1.70)
Table 2: False negatives identification results. We report the average of recall, precision, and f1-score of the identified false negatives over the final epoch of pretraining.
Method Precision Recall F1-Score
SogCLR + FNC 27.57 (0.03) 53.67 (0.27) 36.42 (0.07)
SogCLR + GloFND 48.40 (0.65) 58.81 (0.60) 53.10 (0.31)
Method CIFAR10 CIFAR100 Food101 Caltech101 Cars
SogCLR 82.46 (0.36) 60.20 (0.22) 59.66 (0.20) 77.73 (0.17) 25.99 (0.64)
+ FNC 82.77 (0.36) 60.96 (0.31) 59.82 (0.14) 78.34 (0.92) 26.28 (0.50)
+ GloFND 82.81 (0.23) 61.94 (0.24) 59.87 (0.16) 79.18 (0.52) 27.88 (0.44)
Method DTD Pets Flowers Average
SogCLR 57.80 (0.30) 60.63 (0.34) 76.91 (0.33) 62.67
+ FNC 58.60 (0.29) 61.67 (0.66) 78.77 (0.51) 63.40 (+0.73)
+ GloFND 58.97 (0.96) 63.91 (0.22) 79.89 (0.09) 64.31 (+1.64)

### v2

@@ Abstract
contrastive learning, false negative discovery, self-supervised learning, machine learning
@@ 1 Introduction
In the absence of reliable labels to determine whether a pair of data is positive or negative, many methods resort to instance discrimination. To this end, positive pairs are defined as distinct augmented views of the anchor data, while negative pairs are generated by sampling from the whole dataset excluding the anchor data, irrespective of their semantics (Chen et al., 2020a; Yuan et al., 2022). However, negative pairs produced through this method lack reliability. Specifically, augmented views from images sharing similar semantic meanings are incorrectly deemed negative, leading to their embeddings being pushed apart. This inadvertently encourages the model to discard crucial semantic information. We term these undesirable negative pairs as false negatives (FN). Figure 1 illustrates examples of false negatives encountered during training on ImageNet, where the anchor images are different from their negative samples yet semantically similar.
The presence of false negatives detrimentally impacts the representations learned through contrastive learning (Saunshi et al., 2019), with this effect becoming more pronounced in large-scale datasets featuring numerous semantic concepts (Chen et al., 2022). For instance, during SogCLR (Yuan et al., 2022) pretraining on ImageNet100 (100 classes), approximately 1% of all negative pairs are false negatives. This translates to around 20,000 false negatives per batch with a batch size of 1024, and about 325 with a batch size of 128. The presence of such false negatives during training can significantly degrade the quality of learned representations: a linear classifier trained on top of such representations achieves up to 10% lower accuracy in a semi-supervised setting compared to representations learned without the false negatives.
@@ 2 Related Work
Early efforts in SSL formulated pretext tasks to enable models to learn representations from unlabeled data. Examples include predicting the relative offset between two patches within the same image (Doersch et al., 2016), solving jigsaw puzzles (Noroozi & Favaro, 2017), colorizing grayscale images (Zhang et al., 2016), and unsupervised deep clustering (Caron et al., 2019). However, these methods necessitate carefully crafted pretext tasks, which may not always apply to diverse domains, leading to a lack of generality.
@@ 3 CL with Global False Negative Identification
$\mathcal{L}_{\mathrm{GCL}}(\mathbf{w})=\mathbf{E}_{\mathbf{x}_{i}\in\mathcal{D},A,A^{\prime}\in\mathcal{P}}\left[\ell(\mathbf{w};\mathbf{x}_{i},A,A^{\prime})\right],$ (1)
@@ 3.1 Learning Dynamic Per-Anchor Thresholds
The hyperparameter $\alpha$ allows GloFND to adapt to different definitions of false negatives, which are inherently dependent on the desired level of granularity. For example, in a dataset like ImageNet, consider two classification settings: (1) a coarse-grained task of classifying between cars and animals, and (2) fine-grained task of classifying dog breeds. Two images of a dog from different breeds might be considered a false negative in the coarse-grained case (1), but not in the fine-grained one (2). Consequently, the optimal percentage of false negatives, controlled by $\alpha$, varies with the chosen granularity. By adjusting $\alpha$, GloFND can flexibly align with different levels of semantic resolution, i.e., granularity. Its value can be set based on prior knowledge (e.g., expected rate of false negatives or desired granularity of learned representations) or tuned like other hyperparameters, such as the temperature in contrastive learning.
@@
Instead, we choose an optimization-based approach to automatically learn the per-anchor threshold $\lambda_{i}$ to select the top $\alpha$most similar negative data for the $i$-th anchor. To achieve this, we cast the problem of finding the $(1-\alpha)$-quantile of all similarity scores between $\mathbf{x}_{i}$ and all other samples, i.e., $R_{i}=\{\mathrm{sim}(E_{\mathbf{w}}(\mathbf{x}_{i}),E_{\mathbf{w}}(\mathbf{x}))\mid\mathbf{x}\in\mathcal{S}_{i}^{-}\}$ as the following optimization problem (Ogryczak & Tamir, 2003):
@@
$\min_{\mathbf{w}}\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})=\frac{1}{n}\sum_{\mathbf{x}_{i}\in\mathcal{D}}\mathbf{E}_{A,A^{\prime}}[\ell(\mathbf{w},\lambda_{i};\mathbf{x}_{i})],$ (3)
$\ell(\mathbf{w},\lambda_{i};\mathbf{x}_{i})=-\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{\prime})+\tau\log(|\widetilde{\mathcal{S}}_{i}^{-}|g(\mathbf{w},\lambda_{i};\mathbf{x}_{i},\widetilde{\mathcal{S}}_{i}^{-})),$
@@
Note that $\min_{\mathbf{w}}\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ can be viewed as a stochastic bilevel optimization problem (Ghadimi & Wang, 2018) since the minimization of $\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ involves the solution $\bm{\lambda}$ to a lower level problem in (2). However, the problem in (3) is more challenging that most bilevel problems in the literature (Ghadimi & Wang, 2018; Ji et al., 2021) because the lower-level problem in (2) is non-smooth and non-strongly convex while the upper-level function $\mathcal{L}_{\mathrm{GCL}}(\mathbf{w},\bm{\lambda})$ is non-differentiable to $\bm{\lambda}$. To tackle this challenge, we just ignore the hypergradient of $\bm{\lambda}$ in terms of $\mathbf{w}$, which has been used in model-agnostic meta-learning (Finn et al., 2017).
@@ 3.2 GloFND for Unimodal CL
We propose an efficient algorithm called GloFND for dynamically discovering and eliminating the false negatives in contrastive learning. GloFND can be combined with previous contrastive learning algorithms, e.g., SogCLR (Yuan et al., 2022). In each iteration, SogCLR + GloFND first randomly samples a batch of data $\mathcal{B}\subset\mathcal{D}$ and data augmentations $A,A^{\prime}$. Then, it alternatively executes two steps: (i) updating the thresholds $\lambda_{i},i\in\mathcal{B}$; (ii) removing the identified false negatives from the loss function and updating the parameters $\mathbf{w}$ of the encoding network. Notably, step (i), GloFND ’s computation of $\lambda_{i}$, is independent of the specific contrastive loss used, as it relies solely on the embedding similarity of negative pairs. The contrastive loss only influences how the filtered false negatives are handled during training.
@@ 3.2.1 Updating the threshold $\lambda$
where $\mathbb{I}(\cdot)$ is the indicator function, $\Pi_{[-1,1]}[\cdot]$ denotes the projection onto the interval $[-1,1]$ due to that are similarity scores are in $[-1,1]$, and $\theta$ is the learning rate of $\lambda_{i}$. In this way, we keep track of a threshold $\lambda_{i}$ for detecting global false negatives across the whole dataset for each anchor $\mathbf{x}_{i}\in\mathcal{D}$, while ensuring that computation remains mini-batch-wise.
@@ 3.2.2 Updating Encoder Network $E_{\mathbf{w}}$
3.3 Extension to Bimodal CL
Our approach GloFND can be extended to resolve the global false negative discovery in bimodal CL, e.g., CLIP (Radford et al., 2021). This can be achieved by learning a threshold for each instance for two modalities and following the same general procedure as for unimodal CL. In this section, we provide a brief overview of GloFND’s extension to bimodal CL.
Let $\mathcal{D}=\{(\mathbf{x}_{1},\mathbf{t}_{1}),\dotsc,(\mathbf{x}_{n},\mathbf{t}_{n})\}$ be a set of image-text pairs, and denote the encoder network by $E_{I}$ for images and the encoder network by $E_{T}$ for text parametrized by $\mathbf{w}$. For each $(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}$, the negative dataset for each anchor image $\mathbf{x}_{i}$ is $\mathcal{S}_{I,i}^{-}=\{\mathbf{t}\mid\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\setminus\{\mathbf{x}_{i},t_{i}\}\}$ while the negative dataset for each anchor text $\mathbf{t}_{i}$ is $\mathcal{S}_{T,i}^{-}=\{\mathbf{x}\mid\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\setminus\{\mathbf{x}_{i},t_{i}\}\}$. Let $\mathbf{z}_{I,i}$ be the representation of the $i$-th anchor image $\mathbf{x}_{i}$ and $\mathbf{z}_{T,i}$ be the representation of the $i$-th anchor text $\mathbf{t}_{i}$, and $\lambda_{I,i},\lambda_{T,i}\in[-1,1]$ their respective associated thresholds defined through (2). Then, the problem is formulated as:
$\min_{\mathbf{w}}\frac{1}{n}\sum_{(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}}\mathbf{E}[\ell(\mathbf{w},\lambda_{I,i},\lambda_{T,i};\mathbf{x}_{i},\mathbf{t}_{i})]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\tau\log|\widetilde{\mathcal{S}}_{I,i}^{-}|g_{I}(\mathbf{w},\lambda_{I,i};\mathbf{x}_{i},\widetilde{\mathcal{S}}_{I,i}^{-})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\tau\log|\widetilde{\mathcal{S}}_{T,i}^{-}|g_{T}(\mathbf{w},\lambda_{T,i};\mathbf{t}_{i},\widetilde{\mathcal{S}}_{T,i}^{-}),$
where $\widetilde{\mathcal{S}}_{I,i}^{-}=\{\mathbf{t}\mid\mathbf{t}\in\mathcal{S}_{I,i}^{-},\mathrm{sim}(\mathbf{z}_{I,i},E_{T}(\mathbf{t}))\leq\lambda_{I,i}\}$ and $\widetilde{\mathcal{S}}_{T,i}^{-}=\{\mathbf{x}\mid\mathbf{x}\in\mathcal{S}_{T,i}^{-},\mathrm{sim}(E_{I}(\mathbf{x}),\mathbf{z}_{T,i})\leq\lambda_{T,i}\}$.
Then, we can easily extend the GloFND to the bimodal setting similarly as in the unimodal setting. We refer readers to the Appendix B for more details.
@@ 4 Experiments
Table 1: Linear evaluation results in unimodal semi-supervised scenario. We train the linear classifiers with different percentages of randomly sampled labeled training data and present their top-1 accuracies (%) on the validation set. We include the overall average and, in parentheses, its improvement WRT SogCLR baseline. We also report the average of recall, precision, and f1-score of the identified false negatives over the final epoch of pretraining.
Method Top-1 Accuracy False Negatives Identification
100.0% 10.0% 1.0% 0.1% Average Precision Recall F1-Score
SogCLR 76.55 (0.09) 72.24 (0.14) 62.92 (0.34) 34.94 (0.60) 61.66
+ FNC 77.12 (0.14) 72.89 (0.25) 64.29 (0.34) 36.11 (0.70) 62.60 (+0.94) 27.57 (0.03) 53.67 (0.27) 36.42 (0.07)
+ GloFND 77.59 (0.03) 73.36 (0.15) 65.09 (0.49) 37.38 (0.97) 63.36 (+1.70) 48.40 (0.65) 58.81 (0.60) 53.10 (0.31)
Method CIFAR10 CIFAR100 Food101 Caltech101 Cars DTD Pets Flowers Average
SogCLR 82.46 (0.36) 60.2 (0.22) 59.66 (0.2) 77.73 (0.17) 25.99 (0.64) 57.80 (0.30) 60.63 (0.34) 76.91 (0.33) 62.67
+ FNC 82.77 (0.36) 60.96 (0.31) 59.82 (0.14) 78.34 (0.92) 26.28 (0.50) 58.60 (0.29) 61.67 (0.66) 78.77 (0.51) 63.40 (+0.73)
+ GloFND 82.81 (0.23) 61.94 (0.24) 59.87 (0.16) 79.18 (0.52) 27.88 (0.44) 58.97 (0.96) 63.91 (0.22) 79.89 (0.09) 64.31 (+1.64)
@@
The unimodal experiments are run on a single NVIDIA A30 with 24GB memory size, while the bimodal experiments make use of a multi-node setup with 2 nodes, each with 2 NVIDIA A100 GPUs with 40GB each.
For unimodal and semi-supervised experiments, we use SogCLR (Yuan et al., 2022) and compare with FNC (Huynh et al., 2022). FNC computes the top $k$ for negative data within a mini-batch by utilizing a support set. The support set includes additional views for each image, and the similarity scores are averaged across these views. For a fair comparison, we set $k=\alpha|\mathcal{D}|$ and use a support set of size 1. For bimodal experiments, we compare GloFND with SogCLR and FastCLIP (Wei et al., 2024).
@@ 4.1 Unimodal and Semi-supervised Experiments
Dataset. We run our experiments on ImageNet100 (Wu et al., 2019), a subset of ImageNet with 100 randomly selected classes (about 128k images), and report scores on its official validation split. Additionally, we examine the transfer learning performance on Food-101 (Bossard et al., 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Stanford Cars (Krause et al., 2013), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Oxford-IIIT Pets (Parkhi et al., 2012), Caltech-101 (Li et al., 2022), and Oxford 102 Flowers (Nilsback & Zisserman, 2008).
Experiment Setup. Following previous work (Yuan et al., 2022), we pretrain ResNet-50 (He et al., 2015) with a 2-layer $128\times 128$ projection head on top of the backbone encoder. We pretrain for 200 epochs with a batch size of 128 and the same set of augmentations as in SogCLR. We use LARS optimizer (You et al., 2017) with square root learning rate scaling ($0.075\times sqrt(BatchSize)$) and cosine decay schedule without restart. For SogCLR, we set the temperature ($\tau$) to 0.1 and $\gamma=0.9$. We start using GloFND when we reach 70 epochs. We use $\alpha=0.01$, initialize $\lambda_{i}=1$, and learn it with Adam with a learning rate of $0.05$ ($\beta_{1}=0.9,\beta_{2}=0.98$) during the remaining epochs. For FNC, we set $\alpha=0.01$ and tune the starting epoch in $\{10,30,50,70,90,110,130\}$, choosing the value that achieves the best semi-supervised average performance. More details on hyperparameters can be found in Appendix D.1.
@@
Results. We report false negative identification performance and top-1 accuracies by linear evaluation in Table 1. GloFND achieves significant improvements in false negative identification over FNC, with a 20.83% and 5.14% increase in mean precision and recall, leading to an F1-score of 53.10%, which is 16.68% higher than FNC. Observe that simply removing the false negatives identified by FNC or GloFND improves both the semi-supervised and transfer learning performance of SogCLR. GloFND achieves greater improvements in both scenarios, achieving 1.04%-2.44% improvement in the semi-supervised scenario and an average 1.64% improvement in transfer learning, while increasing the per-epoch computation by only 2% (from 427 s to 435 s). More details can be found in Appendix D.3. Note these improvements are achieved by simply removing the false negatives identified by GloFND from the loss, while a more careful treatment can potentially improve the performance even further.